Nordic Journal of Computing

INSTRUCTIONS TO USE NJC STYLE FILES

KIMMO E. E. RAATIKAINEN University of Helsinki, Department of Computer Science

P.O. Box 26 (Teollisuuskatu 23), FIN-00014 University of Helsinki, Finland

               Kimmo.Raatikainen@cs.Helsinki.FI

Abstract. Each paper must include an abstract that summarizes the paper. Recommended length is at most 150 words. The abstract should not contain any references or displayed equations. L^{A}TEX-environment is abstract.

Key words: List of keywords (4–6) is optional. L^{A}TEX-environment is keywords.

CR Classification: Each paper should provide the CR Classification (see the appropriate issue of Computing Reviews). L^{A}TEX-environment is subject.

1. Introduction

Nordic Journal of Computing encourages the authors to use L^{A}TEX in preparing the manuscripts with the L^{A}TEX style file njcarticle.sty and the BibTEX style file njcarticle.bst. The figures must be either in the L^{A}TEX picture environment or in Encapsuled Postscript.

This document describes how the L^{A}TEX style file njcarticle.sty and the BibTEX style file njcarticle.bst are to be used. The files can be obtained through anonymous ftp, as follows:

%ftp ftp.cs.Helsinki.FI
Name: anonymous
Password: <your email address>
ftp> cd pub/Nordic_Journal_of_Computing/Style_Files
ftp> get njcarticle.sty
ftp> get njcepsmacros.sty
ftp> get njcarticle.bst
ftp> get njcdoc.tex
ftp> get njcdoc.bib
ftp> get njclogo.ps
ftp> quit

The L^{A}TEX-file of an accepted paper, in the final form, should be send through email to njc@cs.Helsinki.FI. If the paper contains Encapsuled Postscript files and/or BibTEX is used send the files separately and an email that explains the files sent.

Received August 1994. Revised September, October 1994.

2. General

The L^{A}TEX style file njcarticle.sty does not accept any options. In \documentstyle-command all options given in brackets are simply omitted. The layout of page should not be changed.

Redundant spaces ought to be minimized by careful arrangement of tables and figures. Read your .log file carefully. There should be no Overfull \hbox (as here) and certainly no visible one (more than 1pt). If necessary, reword the text. The preamble command \draft can be used to produce a visible overfullrulebox in the margin.

The style file provides macros to create running heads. The \authorhead contains the authors’ surnames as “Author”, “Author1, Author2”, and “Author1, Author2, Author3” for up to three authors, and “Author1 et al.” for four or more authors. The \titlehead contains a short form of the title, not more than 30 characters.

The argument of title-command must be written in capitals. When the paper has more than one author, the authors in the argument of authorcommand are separated by usual \and-command or by \AND-command that inserts vertical glue between the blocks of authors’ names, affiliations, etc.

Neither footnotes nor appendices should be used. However, if appendices are really necessary, their place is after Acknowledgements and before References.

3. Sectioning, numbering, etc.

3.1 Sectioning

The following three L^{A}TEX-sectioning commands are available: \section,

\subsection, and \subsubsection.

3.2 Numbering

The numbering of displayed equations, theorems, figures, tables, and other “numbered” environments follows one of two styles: either consecutive in each section, or consecutive through the whole paper (default). If you prefer the first style, the style file provides \twolevelnumbering-command. The argument of the command is a list of environments that are numbered consecutively in each section, e.g.

    \twolevelnumbering{figure,equation,theorem}

Tables should be referred to “Table I”. Equations and figures should be referred to in abbreviated forms: “Eq. (??)” and “Fig. ??”. Use macros \eqnref{eqn-label}, \figref{fig-label}, and \tableref{tab-label} to generate the references.

3.3 Lists of items

The depths of L^{A}TEX-environments itemize and enumerate are restricted to two.

3.4 Spacing before and after environments

Extra space is added at the top of list if the input file has a blank line before any list-making environment. The vertical space after the environment is the same as the one preceding it. The list-making environments are: quote,

quotation, verse, itemize, enumerate, description, center, flushleft,

and flushright, as well as the theorem-like environments.

3.5 Tabbing environment

The style file defines \tabbingstretch to specify the strut to be used in the tabbing environment. The functionality of \tabbingstretch is the same as that of \arraystretch in array and tabular environments.

4. Theorem-like environments

The style file defines the following theorem-like environments:

         theorem    proposition  claim    fact
         lemma      definition   problem  remark
         corollary  conjecture   example  observation

These are “numbered” environments. The style file also defines the corresponding “unnumbered” environments: theorem*, ..., observation*. The proofs can be typed within the environment proof.

For example, the L^{A}TEX source

\begin{theorem}[Chebychev's Inequality]
If $X$ is any random variable, then
\begin{equation}
\Pr[|X|\ge a] \le \mbox{\rm E}(X^2)/a^2\;.
\label{eqn:example}
\end{equation}
\end{theorem}

produces

Theorem 1. (Chebychev’s Inequality) If X is any random variable, then

Pr[|X| ≥ a] ≤ E(X^{2})/a^{2} .  (4.1)

Fig. 5.1: An example.

Proof. If F (x) denotes the distribution function of the random variable X, then

Pr[|X| ≥ a] = ∫_{|x|≥a} d F (x) .

Since in the region of integration |x|/a ≥ 1, it follows that

∫_{|x|≥a} d F (x) ≤ \frac{1}{a^{2}} ∫_{|x|≥a} x^{2} d F (x) .

By extending the integration to all values of x, we merely strengthen the inequality:

∫_{|x|≥a} d F (x) ≤ \frac{1}{a^{2}} ∫_{|x|≥a} x^{2} d F (x) ≤ \frac{1}{a^{2}} ∫ x^{2} d F (x) = E(X^{2})/a^{2} . □

When a proof ends with a displayed equation as above, the box “□” should be at the right end of the formula rather than at the beginning of the next line. In this case, the box must be inserted through \qed at the end of the equation, and a blank line must be left between the closing of the equation and the \end{proof}. (See the source of these instructions.)

5. Figures and tables

Figures and tables are to be inserted in the text nearest their first reference. They should be arranged so as not to cause an excessive amount of blank